A car accelerates uniformly from rest to 25/m/s/s in 10 seconds. Sketch the graph for displacement versus time graph
1
Expert's answer
2019-09-16T09:45:23-0400

The displacement of the car is described by equation


"d=v_it+\frac{at^2}{2}=\frac{at^2}{2}"

The acceleration of the car


"a=\frac{\Delta v}{\Delta t}=\frac{25\:\rm m/s-0\: m/s}{10\:\rm s}=2.5\:\rm m/s^2"

So

"d=\frac{2.5t^2}{2}=1.25t^2"